Im stuck on this...
If there are 32 coins in $2.45, then how many nickels are there?
5n+10d = 245


Answer by rfer(16322)

n+d=32
d=32-n
---------------
.05n+.1(32-n)=2.45
.05n+3.2-.1n=2.45
-.05n=-.75
n=15
d=32-n=17
-----------------
.05*15+.1*17=
.75+1.7=2.45